# Horizon Brightened Acceleration Radiation from Massive Vector Fields

### v1

@@ IX Conclusion
In this work, we extended Scully et al.’s acceleration-radiation framework Scully:2017utk from a scalar field to a massive spin-1 (Proca) field for atoms freely falling into a Schwarzschild background, treating both detector realizations: a charged-monopole current coupling and a physical electric-dipole coupling. For both detector models, the near-horizon, counter-rotating amplitude reduces to the same Laplace integral, giving the universal spectral kernel in (73) and the detailed-balance ratio in (79). This arises from the local Rindler mapping of $t-r_{*}$ and is insensitive to spin, mass, polarization, or detector realization.
@@
Absolute excitation probabilities acquire smooth prefactors and a hard mass threshold via $k_{\infty}(\nu)$, greybody transmission $\Xi_{\ell}(\nu)$, and polarization/geometry weights. For instance, see the general structure (75) and its current/dipole specializations (76)-(77). Near threshold, polar $\ell=0$ turns on most gently, while axial channels are more strongly suppressed.
@@
Converting probabilities into escaping emission/absorption rates yields the detailed-balance relation (87). The single-mode master equation relaxes to the geometric steady state with mean occupancy given in (92). Detector, mass, and polarization only set relaxation speeds and flux magnitudes, not the equilibrium ratio.
@@
The master equation implies the universal mode-by-mode entropy flux rule (100). Combined with energy balance and the Schwarzschild area law, this gives the HBAR-style area-entropy relation (106). The proportionality constant is universal; all Proca specifics enter only through the radiative area change via the net flux.
The Proca spectrum displays a sharp onset at the mass threshold, channel-dependent near-threshold scaling governed by $\Xi_{\ell}(\nu)$ and $k_{\infty}(\nu)$, a high-frequency regime where $\Xi_{\ell}\to 1$ and the Planckian tail dominates, and order-unity polarization weights distinguishing current versus dipole couplings. These are the experimentally relevant imprints.
@@
The large-$\omega$ expansion confines corrections to the prefactor at order $\nu/\omega$, leaving the thermal kernel (73) intact. The Proca constraint $\nabla_{\mu}A^{\mu}=0$ replaces gauge freedom cleanly, and the massless limit is smooth at the level of probabilities/rates. The factorization into a universal kernel times a smooth prefactor justifies the master-equation treatment and the entropy law for massive spin-1 quanta.
@@
For future directions, explicit axial/polar greybody factors, rotating backgrounds (superradiance), alternative exterior states (Unruh/Hartle-Hawking), and detector-engineering for longitudinal/transverse isolation would sharpen phenomenology without altering the core results (73), (92), and (106).

### v2

@@ IX Conclusion
In this work, we extended Scully et al.’s acceleration-radiation framework [44] from a scalar field to a massive spin-1 (Proca) field for atoms freely falling into a Schwarzschild background, treating both detector realizations: a charged-monopole current coupling and a physical electric-dipole coupling. For both detector models, the near-horizon, counterrotating amplitude reduces to the same Laplace integral, giving the universal spectral kernel in (V.13) and the detailed-balance ratio in (V.20). This arises from the local Rindler mapping of $t-r_{*}$ and is insensitive to spin, mass, polarization, or detector realization.
@@
Absolute excitation probabilities acquire smooth prefactors and a hard mass threshold via $k_{\infty}(\nu)$, greybody transmission $\Xi_{\ell}(\nu)$, and polarization/geometry weights. For instance, see the general structure (V.15) and its current/dipole specializations (V.16)–(V.18). Near threshold, polar $\ell=0$ turns on most gently, while axial channels are more strongly suppressed.
@@
Converting probabilities into escaping emission/absorption rates yields the detailed-balance relation (VI.8). The single-mode master equation relaxes to the geometric steady state with mean occupancy given in (VI.13). Detector, mass, and polarization only set relaxation speeds and flux magnitudes, not the equilibrium ratio.
@@
The master equation implies the universal mode-by-mode entropy flux rule (VII.7). Combined with energy balance and the Schwarzschild area law, this gives the HBAR-style area-entropy relation (VII.13). The proportionality constant is universal; all Proca specifics enter only through the radiative area change via the net flux.
The Proca spectrum displays a sharp onset at the mass threshold, channel-dependent near-threshold scaling governed by $\Xi_{\ell}(\nu)$ and $k_{\infty}(\nu)$, a high-frequency regime where $\Xi_{\ell}\to 1$ and the Planckian tail dominates, and polarization-sensitive weights distinguishing current from dipole couplings. Taken together, these imply a characteristic gap-plus-tail spectral profile with longitudinal/transverse and axial/polar discrimination that is absent in the scalar problem. This is the concrete physical payoff of the Proca extension: it identifies the observable spin-1 signatures that survive on top of the same universal near-horizon thermal kernel. We emphasize, however, that the present paper provides the analytic detector-theory template for those signatures rather than a complete astrophysical detectability analysis.
@@
The large-$\omega$ expansion confines corrections to the smooth prefactor, with relative size controlled by $\mathcal{O}((1+|\nu|)/\omega)$ in the cavity-selected regime, leaving the universal thermal kernel (V.13) and the leading detailed-balance structure intact.. The Proca constraint $\nabla_{\mu}A^{\mu}=0$ replaces gauge freedom cleanly, and the massless limit is smooth at the level of probabilities/rates. The factorization into a universal kernel times a smooth prefactor justifies the master-equation treatment and the entropy law for massive spin-1 quanta.
@@
For future directions, explicit axial/polar greybody factors, rotating backgrounds (superradiance), alternative exterior states (Unruh/Hartle-Hawking), and detector-engineering for longitudinal/transverse isolation would sharpen the semirealistic phenomenology developed here. In particular, they would turn the present analytic signatures:threshold gap, channel-dependent turn-on, and polarization-sensitive response, which are into a more quantitative bridge between the HBAR detector framework and massive-vector scenarios of current interest.